OpenStudy (anonymous):

for shipment a women's basketball is packaged in a cubed box. The circumference of the basketball is 28.5in. what is the volume of the smallest cubed box that will hold the basketball. answer to the nearest cubic inch. can you explain how you got the answer to thanks.

OpenStudy (anonymous):

remember the equation for a circle a sphere is just a circle rotated around pi radians so the circumference of a circle is circ = 2*pi*r since d = 2*r where d is the diameter circ = d*pi solving for d d = circ /pi so we have the solution for diameter that means the box that would hold this ball would have to be d x d x d this means v = circ/pi* circ/pi*circ/pi v = circ^3/pi^3 that would be the smallest volume v = 747 cubic inches hope that helps good luck

OpenStudy (anonymous):

that actually did thanks.